The length of a rectangle is 4 times the width, and the area is 256 square feet. Find the length and the widthe of the rectangle.
Answer by ptaylor(2198)


Let x=width
Then 4x=length
Area of a rectangle = Length times Width=4x*x
So:
4x^2=256 divide both sides by 4
x^2=64 take sqrt of both sides
x=+-8 cm discount negative value for x.
x=8 cm--------------width
4x=4*8=32 cm-----------length
CK
A=8*32=256
256=256

Hope this helps----ptaylor